• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ipom

    Log entry time 03:18:09 on June 27, 2010

    Entry number 328633

    keyword=HV changes to s2m

    We used Matt's script of peak fitting of the s2m. We ran the script on run 1207 and updated HV for all channels.






    Figure 1

    HVcalibration_1207.ods